OpenStudy (twizttiez):

What is the volume of a cone that has a radius of 4 cm and a height of 9 cm? Use 3.14 to approximate pi and express your answer in hundredths. cm3

OpenStudy (twizttiez):

@hockeychick23

OpenStudy (hockeychick23):

V= πr^2(h/3) V=π (4^2)(9/3) V=150.8